Before the distribution of gadgets, the teacher will test two tablets. Use D for defective tablet, Nfor
non-defective tablet.
Let X be the random variable representing the number of defective tablet that will occur.
Complete the table below and find the values of random variable X.
1
Expert's answer
2021-09-20T18:08:39-0400
"\def\arraystretch{1.5}
 \begin{array}{c:c}
 \text{Possible } & \text{Values of the } \\
 \text{Outcomes} & \text{Random Variable} \ L\\
\hline
LL & 2 \\
LY & 1 \\
YL & 1\\
YY & 0

\end{array}""P(L=0)=\dfrac{\dbinom{3}{0}\dbinom{10-3}{2-0}}{\dbinom{10}{2}}=\dfrac{21}{45}=\dfrac{7}{15}""P(L=1)=\dfrac{\dbinom{3}{1}\dbinom{10-3}{2-1}}{\dbinom{10}{2}}=\dfrac{21}{45}=\dfrac{7}{15}""P(L=2)=\dfrac{\dbinom{3}{2}\dbinom{10-3}{2-2}}{\dbinom{10}{2}}=\dfrac{3}{45}=\dfrac{1}{15}""\def\arraystretch{1.5}
 \begin{array}{c:c}
 l & 0 & 1 & 2 \\ 
 \hdashline \\

 P(L=l) & \dfrac{7}{15} & \dfrac{7}{15} & \dfrac{1}{15}
\end{array}"